Project S costs $15,000, and its expected cash flows would be $4,500 per year for 5 years. Mutually exclusive Project L costs $37,500, and its expected cash flows would be $11,100 per year for 5 years. If both projects have a WACC of 14%, which project(s) would be accepted?

A)

Neither Project S or Project L would be accepted.

B)

Project S and Project L would be both accepted.

C)

Project L would be accepted.

D)

Project S would be accepted.

E)

The information provided is not enough to make a decision.
Answer #1

Hi

We will find the NPV for both the projects to select the project.

NPV for Project S = -15000 + 4500/(1+14%)^1 + 4500/(1+14%)^2+ 4500/(1+14%)^3+ 4500/(1+14%)^4+ 4500/(1+14%)^5

=-15000 + 3947.37 + 3462.60 + 3037.37 + 2664.36 + 2337.16

=$448.86

NPV for project L = -37500 + 11100/(1+14%)^1 + 11100/(1+14%)^2 + 11100/(1+14%)^3 + 11100/(1+14%)^4 + 11100/(1+14%)^5

=-37500 + 9736.84 + 8541.09 + 7492.18 + 6572.09 + 5764.99

=$607.20

Since both projects are having positive NPV hence option B is correct(Both projects should be selected.)

Thanks